# Search for GeV-PeV neutrinos from nova T Coronae Borealis with IceCube

### v1

@@ 2 Sensitivity and Neutrino Detection Prospects
RS Oph 2021 | 4.5 | 6.47 | 29.50 | |
T CrB 1866 | 2.0 | 20.73 | 7.25 | |
= 10.12 RS Oph | =3.21 RS Oph | |||
T CrB 1946 | 3.0 | 20.71 | 6.25 | |
= 4.03 RS Oph | =3.20 RS Oph |
For the IceCube search we plan a conservative time window, days relative to the optical peak, in order to cover the entire time and possible early or late emission. The sensitivity and discovery potential at this declination as a function of the analysis time window are shown in Figure 4. The short duration of this nova compared to RS Oph is advantageous for neutrino follow-up, as it reduces the expected number of background events in the direction of the nova.

### v2

@@ 2 Sensitivity and Neutrino Detection Prospects
| RS Oph 2021 | 4.5 | 8.87 | 37.50 | |
| T CrB 1866 | 2.0 | 52.32 | 7.75 | |
| = 10.12 RS Oph | =5.90 RS Oph | |||
| T CrB 1946 | 3.0 | 35.09 | 8.25 | |
| = 4.03 RS Oph | =3.95 RS Oph |
For the IceCube search we plan a time window of days relative to the optical peak, in order to cover the time and possible early or late emission. The sensitivity and discovery potential at this declination as a function of the analysis time window are shown in Figure 4. The short duration of this nova compared to RS Oph is advantageous for neutrino follow-up, as it reduces the expected number of background events in the direction of the nova.